OpenStudy (anonymous):

Find the Gcp: 18x^3 and 30x^5

OpenStudy (anonymous):

factor each expression

OpenStudy (anonymous):

18x^3=18*x*x*x=6*3*x*x*x=3*3*2*x*x*x

OpenStudy (anonymous):

30x^5=6*5*x*x*x*x*x=3*2*5*x*x*x*x*x

OpenStudy (anonymous):

the common are x*x*x and 6

OpenStudy (anonymous):

so GCF=6x^3